OpenStudy (anonymous):

The function f(x, y) = 1−x over the region R : [0, 1]×[0, 1] on xy-plane forms a triangular solid. Sketch the solid.

OpenStudy (amistre64):

well, first thing to do is to draw up an xyz space ... and form the rectangle in the xy plane

OpenStudy (amistre64):

you sure your f(x,y) is correct?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

well, you f(x,y) is a plane the divides you region in half as opposed to being a cap over it

OpenStudy (amistre64):

pfft, i think i see it off a little; i did y = 1-x, instead of z=1-x

OpenStudy (amistre64):



OpenStudy (amistre64):

thats better

OpenStudy (anonymous):

woooow, thanks so much......